INTRACAVITY LASER ABSORPTION SPECTROSCOPY OF PLATINUM DIMER IN THE NEAR IN-FRARED

\underline{LEAH C. O’BRIEN}, Department of Chemistry, Southern Illinois University, Edwardsville, IL 62026-1652; JAMES J. O’BRIEN, Department of Chemistry and Biochemistry and Center for Nanoscience, University of Missouri, St Louis, MO 63121-4499.

The 12937cm^{−1} band of Pt dimer has been recorded with rotational resolution using intracavity laser absorption spectroscopy. Based on Pt isotope shifts, this band is identified as the (1,0) band of an Ω = 1 − XΩ = 1 transition. The results of the analysis will be presented and compared with previous work.